OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth, if necessary. x2 + 15 =0 {3.9} {3.9, -3.9} {7.5, -7.5} no real solution

OpenStudy (phi):

is there a typo in this question? Don't you mean x^2 -15=0 ?

OpenStudy (phi):

but if it is \[ x^2 +15=0 \] add -15 to both sides \[ x^2 +15-15 =0-15 \] simplify. what do you get ?